# QVGen: Pushing the Limit of Quantized Video Generative Models

### v4

@@ Abstract
Video diffusion models (DMs) have enabled high-quality video synthesis. Yet, their substantial computational and memory demands pose serious challenges to real-world deployment, even on high-end GPUs. As a commonly adopted solution, quantization has achieved notable successes in reducing cost for image DMs, while its direct application to video DMs remains ineffective. In this paper, we present QVGen, a novel quantization-aware training (QAT) framework tailored for high-performance and inference-efficient video DMs under extremely low-bit quantization (i.e., $4$-bit or below). We begin with a theoretical analysis demonstrating that reducing the gradient norm is essential to facilitate convergence for QAT. To this end, we introduce auxiliary modules ($\Phi$) to mitigate large quantization errors, leading to significantly enhanced convergence. To eliminate the inference overhead of $\Phi$, we propose a rank-decay strategy that progressively eliminates $\Phi$. Specifically, we repeatedly employ singular value decomposition (SVD) and a proposed rank-based regularization $\bm{\gamma}$} to identify and decay low-contributing components. This strategy retains performance while zeroing out additional inference overhead. Extensive experiments across $4$ state-of-the-art (SOTA) video DMs, with parameter sizes ranging from $1.3\text{B}\sim 14\text{B}$, show that QVGen is the first to reach full-precision comparable quality under $4$-bit settings. Moreover, it significantly outperforms existing methods. For instance, our $3$-bit CogVideoX-22B achieves improvements of +25.28+25.28 in Dynamic Degree and +8.43+8.43 in Scene Consistency on VBench.
@@ 1 Introduction
Figure 1: Comparison of samples generated by CogVideoX-22(Yang et al., 2025) with a fixed random seed. “WxxAyy” denotes “xx”-bit per-channel weight and “yy”-bit per-token activation quantization. Our approach far outperforms previous PTQ (i.e., (f)) and QAT (i.e., (c)-(e)) methods. To be noted, methods (c)-(f) have achieved noticeable performance for $4$-bit image DMs. More visual results can be found in Sec. L.
@@ 3 QVGen
Figure 2: Overview of the proposed QVGen. (a) This framework integrates auxiliary modules $\Phi$ to improve training convergence (Sec. 3.1). (b) To maintain performance while eliminating inference overhead induced by $\Phi$, we design a rank-decay schedule that progressively shrinks the entire $\Phi$ to $\varnothing$ through iteratively applying the following two strategies (Sec. 3.2): (i) SVD to identify the low-impact components in $\Phi$; (ii) A rank-based regularization $\bm{\gamma}$} to decay the identified components to $\varnothing$. A detailed procedure can be found in Sec. B.
@@ 3.1 Improving Convergence with Auxiliary Modules $\Phi$
where $T$ signifies the total number of training iterations and $f_{t}(\cdot)$ is the unknown cost function at iteration $t$. Here, ${\bm{\theta}}_{t}$ represents the parameters of the quantized video DM at training step $t$, constrained within a convex compact set $\mathbb{S}^{d}$, while ${\bm{\theta}}^{*}=\arg\,\min_{{\bm{\theta}}\in\mathbb{S}^{d}}\sum^{T}_{t=1}f_{t}({\bm{\theta}})$ is the optimal parameters. In QAT, ${\bm{\theta}}_{t}$ is updated by gradient descent, with the learning rate ${\bm{\eta}}_{t}$ and gradient ${\bm{g}}_{t}$, as:
$\begin{array}[]{ll}{\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-{\bm{\eta}}_{t}{\bm{g}}_{t}.\end{array}$ (8)
@@
A smaller value of $\frac{R(T)}{T}$ implies a closer convergence to the optimum. Thm. 3.1 (with a proof provided in Sec. C) suggests that for a large $T$ (i.e., $\frac{dD_{\infty}^{2}}{2T\eta_{T}^{m}}$ becomes negligible), minimizing $\|{\bm{g}}_{t}\|_{2}$ is critical for improving convergence behavior of QAT.
A lower $\|{\bm{g}}_{t}\|_{2}$ is typically observed in more stable training processes (Takase et al., 2024; Xie et al., 2024). Therefore, to reduce $\|{\bm{g}}_{t}\|_{2}$, we aim to stabilize the QAT process by mitigating aggressive training losses (e.g., loss spikes(Kumar et al., 2025; Li et al., 2024b). Specifically, we introduce a learnable auxiliary module $\Phi$ to enhance each quantized linear layer of a video DM. This trainable module aims to mitigate severe quantization-induced errors during QAT, thereby preventing aggressive training losses. The forward computation of such a $\Phi$-equipped layer becomes:
@@
where $\Phi({\mathcal{Q}}_{b}({\mathbf{X}}))={\mathbf{W}}_{\Phi}{\mathcal{Q}}_{b}({\mathbf{X}})$. Here, ${\mathbf{W}}_{\Phi}$ is initialized before QAT by the weight quantization error, defined as ${\mathbf{W}}-{\mathcal{Q}}_{b}({\mathbf{W}})$. More initialization approaches for ${\mathbf{W}}_{\Phi}$ can be found in Sec. J.2.
@@
To validate the effectiveness of $\Phi$, we conduct experiments for CogVideoX 22(Yang et al., 2025) and Wan 1.31.3(WanTeam et al., 2025). Compared with the previous SOTA QAT method Q-DM (Li et al., 2023b), the proposed approach exhibits consistently lower $\|{\bm{g}}_{t}\|_{2}$ and reduced training loss, as depicted in Fig. 3. This aligns well with both the theoretical and the empirical analysis discussed earlier. Therefore, incorporating $\Phi$ in QAT effectively reduces the gradient norm and leads to better convergence for QAT. In addition, as evidenced by Fig. 3, the substantial performance degradation of Q-DM (e.g., depicted in Fig. 1) for the video generation task could be attributed to its relatively large $\|{\bm{g}}_{t}\|_{2}$. Besides, we provide further analyses of $\|{\bm{g}}_{t}\|_{2}$ in video generation QAT in Sec. G.
@@ 3.2 Progressively Shrinking $\Phi$ via Rank-Decay
To improve QAT while eliminating the inference overhead, we propose to progressively remove $\Phi$ throughout the training process. This allows the model to benefit from $\Phi$ during QAT, while ultimately yielding a standard quantized model (Li et al., 2023b; Esser et al., 2020b) with no extra inference cost. To achieve this goal, a straightforward solution is to decay all parameters of $\Phi$ directly. However, we have noticed that it is ineffective and suboptimal (see Tab. 5). This calls for a fine-grained decay strategy.
@@
Based on the above analysis, we investigate the contribution of fine-grained components in $\Phi$. Specifically, we apply singular value decomposition (SVD) (Li et al., 2025; 2023c) to ${\mathbf{W}}_{\Phi}$ at various training steps:
@@
These findings suggest that an increasing number of orthonormal directions $\{{\bm{u}}_{s},{\bm{v}}_{s}\}$ contribute little, as their associated singular values $\sigma_{s}$ are small (Zhang et al., 2015; Yang et al., 2020). Hence, as training proceeds, only an increasingly low-rank portion of $\Phi$ (i.e., $\Sigma^{d^{\prime}}_{s=1}\sigma_{s}{\bm{u}}_{s}{\bm{v}}_{s}^{\top}$, where $d^{\prime}<d$) is needed, and the remaining components can be decayed without noticeably affecting performance. Motivated by this, we propose a novel rank-decay schedule that progressively shrinks $\Phi$ by repeatedly identifying and eliminating the above-mentioned low-impact parts. First, to attain these parts, we reformulate the computation of $\Phi$ as:
@@ 4.1 Implementation Details
Evaluation. We select $8$ dimensions in VBench (Huang et al., 2024b) with unaugmented prompts to comprehensively evaluate the performance following previous studies (Zhao et al., 2025a; Ren et al., 2024). Moreover, for huge ($\geq$ 55B parameters) DMs, we additionally report the results on VBench-2.0 (Zheng et al., 2025a) with augmented prompts to measure the adherence of videos to physical laws, reasoning, etc. More detailed experimental setups can be found in Sec. D.
Table 1: Performance comparison across different quantization methods on VBench (Huang et al., 2024b). “†” indicates PTQ methods and “*” signifies QAT methods. “Full Prec.” denotes the BF16 model. $\clubsuit$” represents that we apply fine-grained per-group weight-activation quantization with a group size of $64$ and keep some linear layers unquantized, which is the same as the official settings of SVDQuant (Li et al., 2025) (details can be found in Sec. D). Best and second-best results are highlighted in bold and underline formats, respectively.
Method
#Bits
(W/A)
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Motion
Smoothness
$\uparrow$
Dynamic
Degree
$\uparrow$
Background
Consistency
$\uparrow$
Subject
Consistency
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@ 4.2 Performance Analysis
Comparison with baselines. We report VBench score comparisons in Tab. 1. In W4A4 quantization, recent QAT methods (Esser et al., 2020b; He et al., 2024; Li et al., 2023b) show non-negligible performance degradation. With W3A3, performance drops become more pronounced. In contrast, the proposed QVGen achieves substantial performance recovery in $3$-bit models and comparable results to full-precision models in $4$-bit quantization. Specifically, it shows higher scores or less than a $2\%$ decrease in all metrics for W4A4 CogVideoX-22(Yang et al., 2025), except Scene Consistency. For PTQ baselines (Zhao et al., 2025a; Li et al., 2025), all fail to generate meaningful content in W4A4 per-channel and per-token settings. Therefore, we apply W4A6 quantization for these methods and also utilize fine-grained per-group W4A4 quantization for SVDQuant (Li et al., 2025). In these cases, W4A4 QVGen outperforms them by a large margin, particularly with $8.37$ and $14.61$ higher Aesthetic Quality and Subject Consistency for Wan 1.31.3B compared to W4A4 SVDQuant. In addition to these findings, we observe that for Wan 1.31.3B, Dynamic Degree recovers easily during QAT, even surpassing that of the BF16 model. However, for CogVideoX-22B, this metric significantly drops. Moreover, Scene Consistency is the most challenging metric to maintain across models and methods.
(a) CogVideoX1.5-55(Yang et al., 2025)
(b) Wan 1414(WanTeam et al., 2025)
Beyond the quantitative results, we provide qualitative results in Fig. 1 and Sec. L, where QVGen markedly improves visual quality over prior methods. Although a clear gap remains between $3$-bit and $4$-bit outputs, our approach shifts the Pareto frontier toward higher accuracy at a smaller model size than existing techniques (e.g., $3$-bit QVGen for CogVideoX achieves a superior Aesthetic Quality than previous $4$-bit methods in Tab. 1). We view this initial exploration as setting a direction and providing valuable insight for future work on 22‑bit quantization.
Besides, we provide more comparisons with additional metrics in Sec. H and comparisons under relatively higher bit-width in Sec. E. We also apply our approach to image generation in Sec. K.
Table 2: Performance for huge video DMs on VBench. Comparison with baselines can be found in Sec. I.
Method
#Bits
(W/A)
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Motion
Smoothness
$\uparrow$
Dynamic
Degree
$\uparrow$
Background
Consistency
$\uparrow$
Subject
Consistency
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
Results for huge DMs. To demonstrate the scalability of our method, we further test two huge video DMs, including CogVideoX1.5-55B and Wan 1414B, at $720p$ resolution. As illustrated in Tab. 2, our $3$-bit and $4$-bit models follow the same pattern seen in smaller models (Tab. 1). However, $3$-bit quantization incurs much larger drops on demanding metrics such as Scene and Overall Consistency, underscoring the challenge of pushing these larger models to $3$ bits. In Fig. 5, we further assess the models with VBench-2.0; the W4A4 DMs incur only negligible overall performance loss.
@@ 4.3 Ablation Studies
To demonstrate the effect of each design, we employ W4A4 Wan 1.31.3B with VBench (Huang et al., 2024b) for ablation studies. More ablations can be found in Sec. J.
@@
Method
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Dynamic
Degree
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
$\quad\quad\lambda\quad\quad$
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Dynamic
Degree
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
$\quad\quad r\quad\quad$
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Dynamic
Degree
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
Table 6: Results of different decay strategies. Details of each method can be found in Sec. F. “Rank” denotes the rank-decay strategy in this work.
Decay
Strategy
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Dynamic
Degree
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
Analysis of different fine-grained decay strategies. To further demonstrate the effectiveness of our rank-decay strategy, we evaluate alternative decay strategies in Tab. 6. Inspired by network pruning (Han et al., 2015; 2016), we introduce a “Sparse” strategy that progressively prunes the largest values in ${\mathbf{W}}_{\Phi}$ during training. Additionally, motivated by residual quantization (Li et al., 2017), we design a “Res. Q.” strategy, which first quantizes ${\mathbf{W}}_{\Phi}$ into $\times44$-bit tensors with the same shape and then progressively removes them one by one. Among all these methods, the “Rank” strategy outperforms others across all the metrics by a large margin. Additionally, both “Sparse” and “Res. Q.” strategies require at least $1.8\times$ training hours at the same setups compared with our “Rank” approach (see Sec. F).

### v5

@@ Abstract
Video diffusion models (DMs) have enabled high-quality video synthesis. Yet, their substantial computational and memory demands pose serious challenges to real-world deployment, even on high-end GPUs. As a commonly adopted solution, quantization has achieved notable successes in reducing cost for image DMs, while its direct application to video DMs remains ineffective. In this paper, we present QVGen, a novel quantization-aware training (QAT) framework tailored for high-performance and inference-efficient video DMs under extremely low-bit quantization (i.e., $4$-bit or below). We begin with a theoretical analysis demonstrating that reducing the gradient norm is essential to facilitate convergence for QAT. To this end, we introduce auxiliary modules ($\Phi$) to mitigate large quantization errors, leading to significantly enhanced convergence. To eliminate the inference overhead of $\Phi$, we propose a rank-decay strategy that progressively eliminates $\Phi$. Specifically, we repeatedly employ singular value decomposition (SVD) and a proposed rank-based regularization $\bm{\gamma}$} to identify and decay low-contributing components. This strategy retains performance while zeroing out additional inference overhead. Extensive experiments across $4$ state-of-the-art (SOTA) video DMs, with parameter sizes ranging from $1.3\text{B}{\sim}14\text{B}$, show that QVGen is the first to reach full-precision comparable quality under $4$-bit settings. Moreover, it significantly outperforms existing methods. For instance, our $3$-bit CogVideoX-22B achieves improvements of +25.28+25.28 in Dynamic Degree and +8.43+8.43 in Scene Consistency on VBench. Code and models are available at https://github.com/ModelTC/QVGen.
@@ 1 Introduction
Figure 1: Comparison of samples generated by CogVideoX-22(Yang et al., 2025) with a fixed random seed. “WxxAyy” denotes “xx”-bit per-channel weight and “yy”-bit per-token activation quantization. Our approach far outperforms previous PTQ (i.e., (f)) and QAT (i.e., (c)-(e)) methods. To be noted, methods (c)-(f) have achieved noticeable performance for $4$-bit image DMs. More visual results can be found in Sec. P.
@@ 3 QVGen
Figure 2: Overview of the proposed QVGen. (a) This framework integrates auxiliary modules $\Phi$ to improve training convergence (Sec. 3.1). (b) To maintain performance while eliminating inference overhead induced by $\Phi$, we design a rank-decay schedule that progressively shrinks the entire $\Phi$ to $\varnothing$ through iteratively applying the following two strategies (Sec. 3.2): (i) SVD to identify the low-impact components in $\Phi$; (ii) A rank-based regularization $\bm{\gamma}$} to decay the identified components to $\varnothing$. A detailed procedure can be found in Sec. A.
@@ 3.1 Improving Convergence with Auxiliary Modules $\Phi$
where $T$ signifies the total number of training iterations and $f_{t}(\cdot)$ is the unknown cost function at iteration $t$. Here, ${\bm{\theta}}_{t}$ represents the parameters of the quantized video DM at training step $t$, constrained within a convex compact set $\mathbb{S}^{d}$, while ${\bm{\theta}}^{*}=\arg\,\min_{{\bm{\theta}}\in\mathbb{S}^{d}}\sum^{T}_{t=1}f_{t}({\bm{\theta}})$ is the optimal parameters. In QAT, ${\bm{\theta}}_{t}$ is updated by gradient descent, with the learning rate $\eta_{t}$ and gradient ${\bm{g}}_{t}$, as:
$\begin{array}[]{ll}{\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta_{t}{\bm{g}}_{t}.\end{array}$ (8)
@@
A smaller value of $\frac{R(T)}{T}$ implies a closer convergence to the optimum. Thm. 3.1 (with a proof provided in Sec. B) suggests that for a large $T$ (i.e., $\frac{dD_{\infty}^{2}}{2T\eta_{T}^{m}}$ becomes negligible), minimizing $\|{\bm{g}}_{t}\|_{2}$ is critical for improving convergence behavior of QAT. A lower $\|{\bm{g}}_{t}\|_{2}$ is typically observed in more stable training processes (Takase et al., 2024; Xie et al., 2024). Therefore, to reduce $\|{\bm{g}}_{t}\|_{2}$, we aim to stabilize the QAT process by mitigating aggressive training losses (e.g., loss spikes(Kumar et al., 2025; Li et al., 2024b). Specifically, we introduce a learnable auxiliary module $\Phi$ to enhance each quantized linear layer of a video DM. This trainable module aims to mitigate severe quantization-induced errors during QAT, thereby preventing aggressive training losses. The forward computation of such a $\Phi$-equipped layer becomes:
@@
where $\Phi({\mathcal{Q}}_{b}({\mathbf{X}}))={\mathbf{W}}_{\Phi}{\mathcal{Q}}_{b}({\mathbf{X}})$. Here, ${\mathbf{W}}_{\Phi}$ is initialized before QAT by the weight quantization error, defined as ${\mathbf{W}}-{\mathcal{Q}}_{b}({\mathbf{W}})$. More initialization approaches for ${\mathbf{W}}_{\Phi}$ can be found in Sec. L.2.
@@
To validate the effectiveness of $\Phi$, we conduct experiments for CogVideoX 22(Yang et al., 2025) and Wan 1.31.3(WanTeam et al., 2025). Compared with the previous SOTA QAT method Q-DM (Li et al., 2023b), the proposed approach exhibits consistently lower $\|{\bm{g}}_{t}\|_{2}$ and reduced training loss, as depicted in Fig. 3. This aligns well with both the theoretical and the empirical analysis discussed earlier. Therefore, incorporating $\Phi$ in QAT effectively reduces the gradient norm and leads to better convergence for QAT. In addition, as evidenced by Fig. 3, the substantial performance degradation of Q-DM (e.g., depicted in Fig. 1) for the video generation task could be attributed to its relatively large $\|{\bm{g}}_{t}\|_{2}$. Besides, we provide further analyses of $\|{\bm{g}}_{t}\|_{2}$ in video generation QAT in Sec. H.
@@ 3.2 Progressively Shrinking $\Phi$ via Rank-Decay
To improve QAT while eliminating the inference overhead, we propose to progressively remove $\Phi$ throughout the training process. This allows the model to benefit from $\Phi$ during QAT, while ultimately yielding a standard quantized model (Li et al., 2023b; Esser et al., 2020b) with no extra inference cost. To achieve this goal, a straightforward solution is to decay all parameters of $\Phi$ directly. However, we have noticed that it is ineffective and suboptimal (see Tab. 5). This observation calls for a fine-grained decay strategy.
@@
Therefore, we begin to investigate the contribution of fine-grained components in $\Phi$. Specifically, we apply singular value decomposition (SVD) (Li et al., 2025; 2023c) to ${\mathbf{W}}_{\Phi}$ at various training steps:
@@
These findings suggest that an increasing number of orthonormal directions $\{{\bm{u}}_{s},{\bm{v}}_{s}\}$ contribute little, as their associated singular values $\sigma_{s}$ are small (Zhang et al., 2015; Yang et al., 2020). We posit that the main weight $\mathbf{W}$ gradually learns to absorb quantization errors. As training proceeds, $\mathbf{W}_{\Phi}$ collapses to low rank and plays an ever‑diminishing role in error compensation. This, in turn, implies that only an increasingly low-rank portion of $\Phi$ (i.e., $\Sigma^{d^{\prime}}_{s=1}\sigma_{s}{\bm{u}}_{s}{\bm{v}}_{s}^{\top}$, where $d^{\prime}<d$) is needed, and the remaining components can be decayed without noticeably affecting performance. Motivated by this, we propose a novel rank-decay schedule that progressively shrinks $\Phi$ by repeatedly identifying and eliminating the above-mentioned low-impact parts. First, to attain these parts, we reformulate the computation of $\Phi$ as:
@@ 4.1 Implementation Details
Evaluation. We select $8$ dimensions in VBench (Huang et al., 2024c) with unaugmented prompts to comprehensively evaluate the performance following previous studies (Zhao et al., 2025a; Ren et al., 2024). Moreover, for huge ($\geq$ 55B parameters) DMs, we additionally report the results on VBench-2.0 (Zheng et al., 2025a) with augmented prompts to measure the adherence of videos to physical laws, reasoning, etc. More detailed experimental setups can be found in Sec. D.
Table 1: Performance comparison across different quantization methods on VBench (Huang et al., 2024c). “†” indicates PTQ methods and “*” signifies QAT methods. “Full Prec.” denotes the BF16 model. $\clubsuit$” represents that we apply fine-grained per-group weight-activation quantization with a group size of $64$ and keep some linear layers unquantized, which is the same as the official settings of SVDQuant (Li et al., 2025) (details can be found in Sec. D). “Full Fine-tuning” denotes we fine-tune the model with the same data as QVGen. Best and second-best results are highlighted in bold and underline formats, respectively.
Method #Bits (W/A) Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Background Consistency $\uparrow$ Subject Consistency $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@ 4.2 Performance Analysis
Comparison with baselines. We report VBench score comparisons in Tab. 1. In W4A4 quantization, recent QAT methods (Esser et al., 2020b; He et al., 2024; Li et al., 2023b) show non-negligible performance degradation. With W3A3, performance drops become more pronounced. In contrast, the proposed QVGen achieves substantial performance recovery in $3$-bit models and comparable results to full-precision models in $4$-bit quantization. Specifically, it shows higher scores or less than a $2\%$ decrease in all metrics for W4A4 CogVideoX-22(Yang et al., 2025), except Scene Consistency. For PTQ baselines (Zhao et al., 2025a; Li et al., 2025), all fail to generate meaningful content in W4A4 per-channel and per-token settings. Therefore, we apply W4A6 quantization for these methods and also utilize fine-grained per-group W4A4 quantization for SVDQuant (Li et al., 2025). In these cases, W4A4 QVGen outperforms them by a large margin, particularly with $8.37$ and $14.61$ higher Aesthetic Quality and Subject Consistency for Wan 1.31.3B compared to W4A4 SVDQuant. In addition to these findings, we observe that for Wan 1.31.3B, Dynamic Degree recovers easily during QAT, even surpassing that of the BF16 model. However, for CogVideoX-22B, this metric significantly drops. Moreover, Scene Consistency is the most challenging metric to maintain across models and methods. Detailed training loss curves across QAT methods and a trial of combining SVDQuant (Li et al., 2025) and QVGen can be found in Sec. G and Sec. I, respectively.
Beyond the quantitative results, we provide qualitative results in Fig. 1 and Sec. P, where QVGen markedly improves visual quality over prior methods. Although a clear gap remains between $3$-bit and $4$-bit outputs, our approach shifts the Pareto frontier toward higher accuracy at a smaller model size than existing techniques (e.g., $3$-bit QVGen for CogVideoX achieves a superior Aesthetic Quality than previous $4$-bit methods in Tab. 1). We view this initial exploration as setting a direction and providing valuable insight for future work on 22‑bit quantization.
Besides, we provide more comparisons with additional metrics in Sec. J and comparisons under relatively higher bit-width in Sec. E. We also apply our approach to image generation in Sec. M.
Table 2: Performance for huge video DMs on VBench. Comparison with baselines can be found in Sec. K.
Method #Bits (W/A) Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Background Consistency $\uparrow$ Subject Consistency $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
Results for huge DMs. To demonstrate the scalability of our method, we further test two huge video DMs, including CogVideoX1.5-55B and Wan 1414B, at $720p$ resolution. As illustrated in Tab. 2, our $3$-bit and $4$-bit models follow the same pattern seen in smaller models (Tab. 1). However, $3$-bit quantization incurs much larger drops on demanding metrics such as Scene and Overall Consistency,
(a) CogVideoX1.5-55(Yang et al., 2025)
(b) Wan 1414(WanTeam et al., 2025)
underscoring the challenge of pushing these larger models to $3$ bits. In Fig. 5, we further assess the models with VBench-2.0; the W4A4 DMs incur only negligible overall performance loss.
@@ 4.3 Ablation Studies
To demonstrate the effect of each design, we employ W4A4 Wan 1.31.3B with VBench (Huang et al., 2024c) for ablation studies. More ablations can be found in Sec. L.
@@
Method Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
$\quad\quad\lambda\quad\quad$ Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
$\quad\quad r\quad\quad$ Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
Analysis of different fine-grained decay strategies. To further demonstrate the effectiveness of our rank-decay strategy, we evaluate alternative decay strategies in Tab. 6. “Linear” in the table denotes linearly reducing the magnitude of the entire full-rank ${\mathbf{W}}_{\Phi}$ to $\mathbf{0}$ by set $\lambda=1$ with a linear schedule as $u$ for Eq. (13). Inspired by network pruning (Han et al., 2015; 2016), we introduce a “Sparse” strategy that progressively prunes the largest values in ${\mathbf{W}}_{\Phi}$ during training. Additionally, motivated by residual quantization (Li et al., 2017), we design a “Res. Q.” strategy, which first quantizes ${\mathbf{W}}_{\Phi}$ into $\times44$-bit tensors with the same shape and then progressively removes them one by one. Among all these methods, the “Rank” strategy outperforms others across all the metrics by a large margin. Additionally, both “Sparse” and “Res. Q.” strategies require at least $1.8\times$ training hours at the same setups compared with our “Rank” approach.
Table 6: Results of different decay strategies. Details of these methods can be found in Sec. F. “Rank” denotes the rank-decay strategy in this work.
Decay Strategy Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
Linear 60.82 52.81 73.19 13.34 21.87
@@
In addition, we introduce two stronger baselines: “Sparse w/ Wanda” and “Sparse w/ MaskLLM”. Rather than pruning the smallest magnitudes as in “Sparse”, the former employs Wanda (Sun et al., 2024) to prune ${\mathbf{W}}_{\Phi}$ using $128$ randomly selected training samples in each decay phase (see “Sparse” in Sec. F). Following MaskLLM (Fang et al., 2024a), the latter applies learned 2:42{:}4 structured pruning masks to ${\mathbf{W}}_{\Phi}$ in each phase. All other settings match those of “Sparse”. “Sparse w/ Wanda” yields a small improvement over “Sparse”, while “Sparse w/ MaskLLM” provides a larger gain; however, both remain below “Rank” on all metrics. It is also worth noting that “Sparse w/ Wanda” requires $1.8\times$ the training time of “Rank”, similar to “Sparse”, and “Sparse w/ MaskLLM” requires $2.1\times$ due to mask learning.